OpenStudy (liizzyliizz):

for the normal (perpendicular) line to the curve y= square root of 8-x^2 at (-2,2) would the slope be 1/2?

OpenStudy (amistre64):

what is f'(-2) ?

OpenStudy (amistre64):

f(x) = 8-x^2 f'(x) = -2x

OpenStudy (amistre64):

the derivative gives us the tangent slope at any given point

OpenStudy (amistre64):

a perp slope (the normal) is the negative reciprocal

OpenStudy (amistre64):

\[-\frac{1}{-2x}=\text{normal slope}\] when x=-2 we get -1/4 if i did it right

OpenStudy (liizzyliizz):

well the choices are -2, 1/2 , -1/2 , 1 , -1 i thought it was 1/2 because when i did the work i ended it up with what u had but i thought the negative signs would cancel eachother out.

OpenStudy (amistre64):

3 negatives is still negative :/

OpenStudy (amistre64):

opps, i didnt read the square root of part

OpenStudy (liizzyliizz):

i didnt see the third negative grrrrr. -_____- i was so close ! :[

OpenStudy (liizzyliizz):

lol and to think i thought i did this right i felt proud for a second.

OpenStudy (amistre64):

sqrt(8-x^2) is easier to see mathically ;)

OpenStudy (amistre64):

you did good, those signs still fool me too

OpenStudy (liizzyliizz):

hmm so it is -1/2 ? aww I was so close.

OpenStudy (amistre64):

\[[\sqrt{8-x^2}]'=\frac{-2x}{2\sqrt{8-x^2}}\] \[-\frac{\sqrt{8-x^2}}{-x}\] \[-\frac{\sqrt{8-(-2)^2}}{-(-2)}\] \[-\frac{\sqrt{8-4}}{2}\] \[-\frac{2}{2}=-1\] so did I do it right?

OpenStudy (amistre64):

it might be easier to solve for the f' and then flip and negate it: \[\frac{-x}{\sqrt{8-x^2}}\] \[\frac{-(-2)}{\sqrt{8-(-2)^2}}\] \[\frac{2}{\sqrt{8-4}}\] \[\frac{2}{\sqrt{4}}=\frac{2}{2}=1\] flip and negate to -1

OpenStudy (liizzyliizz):

hmmm I see what you did, ehh calculus <.< I feel like a scrub now lol.

OpenStudy (amistre64):

;) what were you trying?

OpenStudy (liizzyliizz):

I started in the right the right direction, but after finding the derivative idk what I did to be honest lol.

OpenStudy (amistre64):

keep at it, youll do fine, good luck!

OpenStudy (liizzyliizz):

Haha thanks :DD